(5, l)is a solution to the equation 2x-3y=7. What does this mean?
Answer by stanbon(75887)

(5, l)is a solution to the equation 2x-3y=7. What does this mean?
---
If you put 5 in place of "x" and you put 1 in place of "y"
you will get a true statement:
2*5-3*1 = 10-3 = 7
=======================
Cheers,
Stan H.